Using the data Table 7A calculate the total average daily inpatient census, total average length of stay for newborns, total inpatient bed occupancy rate, and the gross death rate. Show our work.

Report 1: Essex General Hospital for June

Bed Count                                           300

Adults and Children                        275

Newborn                                             25

_______________________________

Admissions                                         1,035

Adults and Children                           894

Newborn                                                141

________________________________

Discharges (including deaths)     1,003

Adults and Children                           865

Newborn                                                138

________________________________

Inpatient Service Days                   9,027

Adults and Children                        8,350

Newborn                                             756

_______________________________

Total Length of Stay                        8,902

Adults and Children                        8,350

Newborn                                                552

______________________________

Deaths                                                   32

<48 hours                                            11

>48 hours                                            21

Adults and Children                         29

Newborn                                                 3

Fetal                                                          8
Answer #1

Admission - 1035

Discharge - 1003

Total average daily inpatient census - 1035 - 1003= 32

Total length of stay 8,902

Adults and children 8350

Total Average length stay of newborn = 8,902 - 8350 = 552

Total inpatient bed occupancy rate

Adults and children = 275

New born = 25

Total inpatient bed occupancy rate 275+25= 300

Gross death rates:

< 48 hours - 11

> 48hours -21

Adults and children -29

Newborn -3

Fetal - 8

Gross death rate total 72